# Charged black hole solutions in $f(R,T)$ gravity coupled to nonlinear electrodynamics

### v1

@@ I Introduction
As GR encounters limitations in explaining certain phenomena, such as dark matter, cosmic inflation, and the late-time acceleration of the universe (attributed to dark energy), alternative approaches have emerged to deepen our understanding of these mysteries. One prominent avenue involves generalizations of GR that replace the Einstein-Hilbert action with arbitrary functions of the Ricci scalar $R$, leading to the so-called $f(R)$ modified theories of gravity [6, 7, 8, 9]. A notable example is Starobinsky’s theory, which incorporates quantum corrections to inflationary models by introducing an $R^{2}$ term into the Einstein-Hilbert action [10]. Another significant extension is the inclusion of the trace of the energy-momentum tensor, $T$, in the gravitational action, giving rise to $f(R,T)$ gravity [11, 12]. Here, $T$ represents the trace of the energy-momentum tensor $T_{\mu\nu}$. The presence of explicit nonminimal curvature-matter couplings result in a nonzero covariant derivative of the energy-momentum tensor [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. This leads to deviations from geodesic motion and the emergence of an extra force [13, 14]. Since its inception, $f(R,T)$ gravity has been applied across a wide range of contexts. These include studies in thermodynamics [23], analyses of energy conditions [24, 25], investigations of compact stars [26, 27, 28, 29], and gravastar solutions [30, 31]. Moreover, it has been extensively explored in cosmology [32, 33, 34] and in the context of wormhole solutions [35, 36, 37, 38, 39]. These alternative frameworks not only extend the scope of GR but also offer promising pathways to addressing some of the universe’s most profound and unresolved questions.
@@
Accordingly, the primary objective of this manuscript is to investigate black hole solutions within the framework of $f(R,T)$ gravity, incorporating nonlinear electrodynamics (NLED) as the matter source. To achieve this, we adopt a model characterized by the functional form $f(R,T)=R+\beta T$, where $\beta$ is a constant. Notably, when $\beta=0$, this model reduces to GR, thereby serving as a generalization of Einstein’s theory. In addition to the gravitational framework, we consider a Lagrangian for the NLED field expressed as a power-law form: $\mathcal{L}=f_{0}+F+\alpha F^{p}$, where $F$ is the Maxwell invariant, $\alpha$ is a coupling constant, and $p$ represents the power-law index. The constant term $f_{0}$ ensures the correct asymptotic behavior. Importantly, setting $\alpha=0$ recovers the linear Maxwell Lagrangian, aligning this formulation with standard electromagnetic theory. With these assumptions in place, we derive the black hole metric function by solving the equations of motion. This involves determining explicit solutions for the metric function under specific values of the power-law index $p$ and analyzing the properties of these solutions. A critical aspect of our analysis is to verify the regularity of the black hole solutions by evaluating curvature invariants, such as the Kretschmann scalar, to identify and characterize singularities.
@@ V.1 Outline of the main results
Finally, we investigated the behavior of the black hole shadow predicted by our model and compared it with the shadow size of the supermassive black hole at the center of our galaxy, Sagittarius A* (Sgr A*), as observed by the Event Horizon Telescope (EHT). The goal of this analysis was to constrain the parameter $\beta$ in our model, which plays a crucial role in shaping the shadow’s characteristics. Specifically, we examined the influence of $\beta$ on the shadow radius for three different values of the power $p$: $p=2$, $p=4$, and $p=6$. In the scenarios we analyzed, we observed that the parameter $\beta$ takes large values, and we determined the ranges for the shadow radius of our black hole, which were constrained by the EHT observations of Sgr A*. The estimated intervals for $\beta$ corresponding to each value of $p$ are as follows: $0\leq\beta_{p=2}\lesssim 1.6\times 10^{4}$, $0\leq\beta_{p=4}\lesssim 3\times 10^{10}$, and $0\leq\beta_{p=6}\lesssim 7.8\times 10^{16}$. These intervals progressively increase as the power $p$ increases, highlighting a direct relationship between the value of $p$ and the allowable range for $\beta$. Moreover, we observed that as $\beta$ increases, the radius of the black hole’s shadow also increases for all values of $p$. This indicates that the shadow’s size is highly sensitive to the value of $\beta$, which is consistent across all considered powers.

### v2

@@ I Introduction
As GR encounters limitations in explaining certain phenomena, such as dark matter, cosmic inflation, and the late-time acceleration of the universe (attributed to dark energy), alternative approaches have emerged to deepen our understanding of these mysteries. One prominent avenue involves generalizations of GR that replace the Einstein-Hilbert action with arbitrary functions of the Ricci scalar $R$, leading to the so-called $f(R)$ modified theories of gravity [6, 7, 8, 9]. A notable example is Starobinsky’s theory, which incorporates quantum corrections to inflationary models by introducing an $R^{2}$ term into the Einstein-Hilbert action [10]. Another significant extension is the inclusion of the trace of the energy-momentum tensor, $T$, in the gravitational action, giving rise to $f(R,T)$ gravity [11, 12]. Here, $T$ represents the trace of the energy-momentum tensor $T_{\mu\nu}$. The presence of explicit nonminimal curvature-matter couplings result in a nonzero covariant derivative of the energy-momentum tensor [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. This leads to deviations from geodesic motion and the emergence of an extra force [13, 14]. Since its inception, $f(R,T)$ gravity has been applied across a wide range of contexts. These include studies in thermodynamics [23], analyses of energy conditions [24, 25], investigations of compact stars [26, 27, 28, 29], and gravastar solutions [30, 31]. Moreover, it has been extensively explored in cosmology [32, 33, 34, 35] and in the context of wormhole solutions [36, 37, 38, 39, 40]. These alternative frameworks not only extend the scope of GR but also offer promising pathways to addressing some of the universe’s most profound and unresolved questions.
@@
Beyond the previous discussion, another fundamental limitation of GR arises from one of its most well-known predictions. Karl Schwarzschild formulated a model for a compact object described solely by its mass. However, his solution features a central singularity, which is a region where the laws of physics break down, leading to infinite curvature and the loss of predictive power. Such singularities pose a major theoretical challenge, as they indicate an incomplete description of gravitational phenomena at extreme conditions. To circumvent these regions where spacetime curvature diverges, various extensions of GR have been proposed, aiming to provide a more complete and physically meaningful framework. One promising approach involves coupling Einstein’s equations with electromagnetism. The resulting solutions, known as regular black holes (RBHs), prevent the formation of a central Schwarzschild singularity by modifying the structure of spacetime at small scales. The first RBH model was introduced by Bardeen [102], but it was only with the work of Ayón-Beato and García that this model was derived as an exact solution of GR, provided that the matter source is described by nonlinear electrodynamics (NLED) [103]. NLED extends classical Maxwell electrodynamics by incorporating nonlinear effects, which become significant in strong electromagnetic fields, thereby offering a natural mechanism to regularize singularities. Inspired by these developments, numerous RBH models have been proposed, incorporating NLED as a matter source within the mathematical framework of GR [104, 105, 106, 107, 108]. In a cosmological context, NLED has also been explored within GR, revealing that the accelerated expansion of the universe can naturally emerge from the electromagnetic quantities involved in these solutions, providing an intriguing alternative to dark energy models [109, 110].
Accordingly, the primary objective of this manuscript is to investigate black hole solutions within the framework of $f(R,T)$ gravity, incorporating NLED as the matter source [111]. To achieve this, we adopt a model characterized by the functional form $f(R,T)=R+\beta T$, where $\beta$ is a constant. Notably, when $\beta=0$, this model reduces to GR, thereby serving as a generalization of Einstein’s theory. In addition to the gravitational framework, we consider a Lagrangian for the NLED field expressed as a power-law form: $\mathcal{L}=f_{0}+F+\alpha F^{p}$, where $F$ is the Maxwell invariant, $\alpha$ is a coupling constant, and $p$ represents the power-law index. The constant term $f_{0}$ ensures the correct asymptotic behavior. Importantly, setting $\alpha=0$ recovers the linear Maxwell Lagrangian, aligning this formulation with standard electromagnetic theory. With these assumptions in place, we derive the black hole metric function by solving the equations of motion. This involves determining explicit solutions for the metric function under specific values of the power-law index $p$ and analyzing the properties of these solutions. A critical aspect of our analysis is to verify the regularity of the black hole solutions by evaluating curvature invariants, such as the Kretschmann scalar, to identify and characterize singularities.
@@ IV.2.1 Constraining the $\beta$ parameter
IV.2.2 Constraining the $\alpha$ parameter
Similar to the procedure developed in the previous section, we now constrain the parameter $\alpha$ based on the data obtained by the Event Horizon Telescope (EHT) for Sgr A*, as illustrated in Fig. 8. We analyze the behavior of the shadow radius $r_{sh}$ for the metric function of our model described by Eq. (22), considering the same values for the constants $M$, $q$, $f_{0}$, $f_{1}$, and $\Lambda$ that were used in the constraint of the parameter $\beta$. We investigate three different scenarios corresponding to the power values $p=\{2,4,6\}$.
Regarding the parameter $\beta$, we use the values obtained from the analysis in the previous section. More specifically, for each value of $p=\{2,4,6\}$, we assume the corresponding values of $\beta$: $\beta=10^{3}$, $\beta=10^{10}$, and $\beta=10^{16}$, respectively. Based on this, we compute the shadow radius using Eq. (59) from the models given by Eqs. (28), (34), and (40). The size of the shadow of the black hole Sgr A* is then represented as a function of the parameter $\alpha$ in Fig. 8, where three different curves are shown: red, yellow, and blue, corresponding to $p=2$, $p=4$, and $p=6$, respectively.
For the case where $p=2$ and $\beta=10^{3}$, represented by the red curve, the shadow radius begins to exceed the limits set by EHT observations when the parameter $\alpha$ approaches the value 7. For smaller values of $\alpha$, the shadow radius remains within the allowed range, as illustrated in Fig. 8.
For $p=4$ and $\beta=10^{10}$, the shadow radius starts to deviate from the constraints imposed by the estimated value for Sgr A* when $\alpha\sim 1.4$. However, for values of $\alpha$ smaller than this, the shadow radius of our model remains consistent with the EHT observational limits for Sgr A*, as indicated by the yellow curve in Fig. 8.
Finally, for $p=6$ and $\beta=10^{16}$, the shadow radius begins to deviate from the expected value for Sgr A* at approximately $\alpha\sim 3.8$. For values of $\alpha$ below this limit, as shown in Fig. 8, the shadow radius remains in agreement with the EHT observations.
Furthermore, Fig. 8 shows that increasing the parameter $\alpha$ leads to an increase in the shadow radius, a trend similar to the behavior observed when constraining the parameter $\beta$. This pattern is maintained across different values of $p$.
Figure 8: The shadow radius $r_{sh}$ for our black hole model (solid red, yellow and blue curve), described by the metric function given in Eq.(22) with power $p=2,4,6$, was calculated according to Eq.(59) for an observer located at $r_{O}\sim 8000\ \text{kpc}$. The shaded regions in blue and light blue represent the EHT constraints for the shadow radius of Sgr A* at confidence levels of $1\sigma$ and $2\sigma$, respectively. The gray regions, beyond the $2\sigma$ values, are excluded by the EHT observations. The dark blue and light blue areas represent regions consistent with the EHT observations of Sgr A*, corresponding to confidence levels of $1\sigma$ and $2\sigma$, respectively.
IV.2.3 Constraining the $p$ parameter
Finally, as in the two previous cases, we constrain the parameter $p$. To achieve this, we use the same values for the constants and set $\alpha=0.5$ and $\beta=10^{3}$. The obtained results provide strong justification for the values of $p$ that were assumed in the development of our solutions. When analyzing the behavior of the shadow radius of our model within the range $p\in[2,6]$, we clearly observe that its values remain entirely within the range compatible with the EHT observations for the black hole Sgr A*. This result validates our choices in constructing the solutions presented in Section III. Fig. 9 illustrates this analysis.
Figure 9: The shadow radius $r_{sh}$ for our black hole model (solid red, yellow and blue curve), described by the metric function given in Eq.(22), was calculated according to Eq.(59) for an observer located at $r_{O}\sim 8000\ \text{kpc}$. The shaded regions in blue and light blue represent the EHT constraints for the shadow radius of Sgr A* at confidence levels of $1\sigma$ and $2\sigma$, respectively. The gray regions, beyond the $2\sigma$ values, are excluded by the EHT observations. The dark blue and light blue areas represent regions consistent with the EHT observations of Sgr A*, corresponding to confidence levels of $1\sigma$ and $2\sigma$, respectively.
@@ V.1 Outline of the main results
Finally, we investigated the behavior of the black hole shadow predicted by our model and compared it with the shadow size of the supermassive black hole at the center of our galaxy, Sagittarius A* (Sgr A*), as observed by the Event Horizon Telescope (EHT). The goal of this analysis was to constrain the parameter $\beta$, $\alpha$ and $p$ in our model, which plays a crucial role in shaping the shadow’s characteristics. Specifically, we examined the influence of $\beta$ on the shadow radius for three different values of the power $p$: $p=2$, $p=4$, and $p=6$. In the scenarios we analyzed, we observed that the parameter $\beta$ takes large values, and we determined the ranges for the shadow radius of our black hole, which were constrained by the EHT observations of Sgr A*. The estimated intervals for $\beta$ corresponding to each value of $p$ are as follows: $0\leq\beta_{p=2}\lesssim 1.6\times 10^{4}$, $0\leq\beta_{p=4}\lesssim 3\times 10^{10}$, and $0\leq\beta_{p=6}\lesssim 7.8\times 10^{16}$. These intervals progressively increase as the power $p$ increases, highlighting a direct relationship between the value of $p$ and the allowable range for $\beta$. Moreover, we observed that as $\beta$ increases, the radius of the black hole’s shadow also increases for all values of $p$. This indicates that the shadow’s size is highly sensitive to the value of $\beta$, which is consistent across all considered powers.
Similar to the procedure that was developed to constrain the $\beta$ parameter, we also constrain and analyze the impact of the $\alpha$ parameter on the behavior of the black hole’s shadow radius for three different values of the power $p=(2,4,6)$. For the $\beta$ parameter, we use the values that are within the margin of constraint we obtained. More specifically, we use: $\beta_{p=2}=10^{3}$, $\beta_{p=4}=10^{10}$, $\beta_{p=6}=10^{16}$. In this context, we note that the intervals for the radius of the shadow of our black hole, limited by the EHT observations of Sgr A*, correspond to the following values of $\alpha$ for each value of $p$: $0\leq\alpha_{p=2}\lesssim 7$, $0\leq\alpha_{p=4}\lesssim 1.4$, $0\leq\alpha_{p=6}\lesssim 3.8$. These intervals show that, as $p$ increases, the value of $\alpha$ decreases. More specifically, the $\alpha$ parameter reaches its lowest value, within observational limits, when $p=4$, and then increases again as $p$ approaches 6. In addition, we observe that as the $\alpha$ parameter increases, the radius of the black hole’s shadow also increases, for all the values of $p$ analyzed.
Finally, to justify the choice of values for $p=(2,4,6)$ used in this manuscript, we also analyzed the influence of the parameter $p$ on the shadow radius of our model. To do this, we used the values $\beta=10^{3}$ and $\alpha=0.5$. As a result, we found that by restricting the parameter $p$, the shadow radius remains completely within the values corresponding to the EHT observations for Sgr A*.